ganeshie8 (ganeshie8):

A triangle is formed by picking \(3\) points at random from the vertices of a regular \(2n+1\)-gon. Find the probability for it to be an equilateral triangle.

ganeshie8 (ganeshie8):

for first few values of \(n\) im getting below probabilities : n = 1, probability = 1 n = 2, probability = 0 n = 3, probability = 0

OpenStudy (thomas5267):

My guess would be for 3n-gon, \(P=\dfrac{n}{\binom{n}{3}}\) and P=0 for all other regular polygon.

ganeshie8 (ganeshie8):

Thats right! could you share how you arrived at that

ganeshie8 (ganeshie8):

pretty sure you meant \(P=\dfrac{n}{\binom{\color{red}{3}n}{3}}\)

OpenStudy (thomas5267):

I am absolutely confused now lol. I thought the answer would be \(\dfrac{3n}{\binom{3n}{3}}\). For all regular polygon of not having 3n vertices, rotation of 120° would not constitute a valid symmetry. So that restricts us to 3n-gon. For 3n-gon, there are \(\binom{3n}{3}\) choices of 3 vertices. Now I have to check my numerator!

OpenStudy (thomas5267):

Ah yes, \(\dfrac{n}{\binom{3n}{3}}\) is indeed the correct answer as there are n equilateral triangle in 3n-gon.

ganeshie8 (ganeshie8):

Looks nice, there are just some repetitions to be accounted for..